OpenStudy (jenniferjuice):

When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?" 1 4 −10 −6

OpenStudy (anonymous):

distribute! 4x(x-3) + 2(x-3)

OpenStudy (anonymous):

Distribute it first\[(a + b)(c + d) = ac + ad + bc + bd\]

OpenStudy (jenniferjuice):

4x^2-12+2x-6 4x^2+2x-18

OpenStudy (anonymous):

No, it's - 12x

OpenStudy (jenniferjuice):

so, 4x^2+2x-12 ?

OpenStudy (anonymous):

No, what I meant is 4x^2 - 12x + 2x - 6 = 4x^2 - 10x - 6

OpenStudy (jenniferjuice):

oh okay so now what do i do? :/

OpenStudy (anonymous):

The standard form of a quadratic equation is y =ax^2 + bx + c. So find b

OpenStudy (jenniferjuice):

do i subtract?

OpenStudy (anonymous):

No, you compare the two equations. y = 4x^2 - 10x - 6 and y = ax^2 + bx + c. What is b?

OpenStudy (jenniferjuice):

-10

OpenStudy (anonymous):

Yes